# SLIM: Scalable and Lightweight LiDAR Mapping in Urban Environments

### v1

@@ VI-D Scalability for Long-term Mapping
Figure 14: Quantitative results of using and without using NFR on the HeLiPR dataset. If the NFR is applied to multi-session data, the time cost and map consumption are kept within certain ranges. Specifically, the BA and NFR take less than 10 seconds. The lightweight map consumes less than 1 MB for one region.
@@
Figure 15: Quantatitive results of using and without using NFR on mapping accuracy, i.e., the absolute trajectory error. The results indicate that the NFR module improves long-term mapping in terms of mapping accuracy.
((a)) Roundabout Region 00
((b)) Roundabout Region 05
Figure 16: The reduced sparse pose graph after applying NFR on HeLiPR. The multi-session data results in multiple pose graphs in the same region. The NFR could sparsity the keyframe and generate a more compact topological map for long-term use.
@@ VI-E Map Re-use for Online Localization
SLIM aims to provide maps that can be reused for online state estimation, i.e., robot localization, a basic need for modern mobile robots. In this section, we employ a straightforward LiDAR-based localization algorithm to evaluate whether our generated map can be utilized for robot relocalization. For simplicity, we primarily assess continuous localization capability, while global relocalization can be achieved using the module presented in Section IV-B2. We apply the same feature extraction operations described in Section IV-A1 to the original LiDAR scans to obtain features as measurements. Subsequently, we use the pose from the previous frame $\mathbf{T}^{w}_{f}$ as the prior pose for the current frame and perform nearest neighbor data association between the points $\mathbf{p}_{l}$ and landmarks $(\mathbf{n}_{k},\mathbf{p}_{k})$ of the current frame, thus obtaining the matching pairs $\mathcal{C}^{\mathcal{L}}$ and $\mathcal{C}^{\mathcal{S}}$. Essentially, localization on the map is to minimize the error between observed points and the pre-built maps. We adopt a point-to-landmarks residual to achieve online localization, described as follows:
@@
Table III reports quantitative results on the HeLiPR dataset. The localization error is generally at the centimeter level, which is sufficiently precise for localization-oriented applications, such as autonomous driving in urban environments. Note that the error includes the errors of pre-mapping, resulting in larger errors in certain regions. For practical applications, additional odometry or GPS information can improve localization accuracy under a sensor fusion scheme. Moreover, the average latency for solving the scan-to-map optimization using a single thread is approximately 42ms (CPU: Intel i7-14650HX), which is sufficient for real-time localization applications.
TABLE III: Online Localization on SLIM-generated Maps: Absolute Trajectory Error (RMSE, Meters)
@@ VII Discussions
Though the designed SLIM system is versatile for long-term LiDAR mapping, potential limitations exist for practical applications. The two key potential limitations are listed as follows:
@@
    A closed-form solution for the NFR requires that the dimension of the recovered residual matches the dimension of the state variables. This means the Jacobian matrix must be square. In general SLAM backends, this condition is not always met. For example, incorporating GPS constraints, pitch-roll constraints, or other constraints may result in mismatched dimensions between residuals and state variables. A crucial factor in our system is that the dimension of lines and planes precisely matches their residual dimension, easily satisfying the condition for a closed-form solution. However, if more information is fused in SLIM, this condition may not hold.
@@
Despite the focus on urban environments in this study, the SLIM system could also be applied to other structured environments, such as forests and built indoor spaces. In forests, the ground and trees can be parameterized into planes and lines. Indoor environments are highly structured with elements like walls and columns, which are typical planes and lines for robot localization. The back-end map merge and refinement can use the same front-end representations for these environments.
@@ VIII Conclusion and Future Study
In this study, we present SLIM, a lightweight and scalable LiDAR mapping system for urban environments. Our key contributions span from front-end feature extraction to back-end map optimization and management. Specifically, we parameterize raw LiDAR point clouds into lines and planes, then design map merge and refinement modules based on these vectorized representations. Experiments conducted on various datasets with different LiDAR sensor types demonstrate the effectiveness of the SLIM system. The results show that SLIM can merge, refine, and maintain new maps with compact representations, creating a scalable mapping framework for long-term use.
Future studies could explore several promising directions. First, SLIM could be deployed in other structured environments similar to urban scenarios. Additionally, at the front end, learning-based feature extraction could reduce the need for parameter tuning. Fusing other sensing modalities, such as visual detection, could enhance landmark stability. At the back end, experiments currently rely on manual region divisions, and a more adaptive approach is needed for long-term autonomy.

### v2

@@ VI-D Scalability for Long-term Mapping
Figure 15: Quantitative results of using and without using marginalization on HeLiPR. The blue solid line represents “Full Map Consumption”; the blue dashed line is “Full Map Consumption (w/o Marginalization)”; the green solid line represents “BA+Marginalization Time” and the green dashed line is “BA Time (w/o Marginalization)”. If the marginalization is enabled, the time cost and map consumption are bounded. More specifically, the BA and marginalization cost less than 10 seconds. The map consumes less than 1 MB for one region.
@@
Figure 16: Quantatitive results of using and without using NFR on mapping accuracy, i.e., the absolute trajectory error. The results indicate that our proposed map-entric NFR performs better than using the prior information matrix, thus demonstrating its effectiveness for the marginalization of long-term mapping.
Figure 17: The reduced sparse pose graph after applying marginalization on HeLiPR. The multi-session data results in multiple pose graphs in the same region. The marginalization provides not only a more compact topological structure but also the information matrix for long-term mapping.
@@ VI-E Map Re-use for Online Localization
SLIM is designed to provide maps that can be reused for online state estimation, i.e., robot localization, a basic need for modern mobile robots. We employ a straightforward LiDAR-based localization algorithm to evaluate whether our generated map can be utilized for robot relocalization. For simplicity, we primarily assess continuous localization capability, while global relocalization can be achieved using the module presented in Section IV-B2. We apply the same feature extraction operations described in Section IV-A1 to the original LiDAR scans to obtain features as measurements. Subsequently, we use the pose from the previous frame $\mathbf{T}^{w}_{f}$ as the prior pose for the current frame and perform nearest neighbor data association between the points $\mathbf{p}_{l}$ and landmarks $(\mathbf{n}_{k},\mathbf{p}_{k})$ of the current frame, thus obtaining the matching pairs $\mathcal{C}^{\mathcal{L}}$ and $\mathcal{C}^{\mathcal{S}}$. Essentially, localization on the map is to minimize the error between observed points and the pre-built maps. We adopt a point-to-landmarks residual to achieve online localization, described as follows:
@@
Table III reports quantitative results on the HeLiPR dataset. The localization error is generally at the centimeter level, which is sufficiently precise for localization-oriented applications, such as autonomous driving in urban environments. Note that the error includes the errors of pre-mapping, resulting in larger errors in certain regions. For practical applications, additional information, such as odometry and semantics [8], could improve localization accuracy under a sensor fusion scheme. Moreover, the average latency for solving the scan-to-map optimization using a single thread is approximately 42ms (CPU: Intel i7-14650HX), which is sufficient for real-time localization applications.
TABLE III: Online Localization on SLIM-generated Maps: Absolute Trajectory Error (RMSE, Centimeters)
@@ VII Discussions
Though the designed SLIM system is versatile for long-term LiDAR mapping, potential limitations exist for practical applications. The two key limitations are listed as follows:
@@
    A closed-form solution for the NFR requires that the dimension of the recovered residual matches the dimension of the state variables. This requires the Jacobian matrix must be square. In SLAM backends, this condition is not always met. For example, incorporating GPS constraints, pitch-roll constraints, or other constraints may result in mismatched dimensions between residuals and state variables. A crucial factor in our system is that the dimension of lines and planes precisely matches their residual dimension, easily satisfying the condition for a closed-form solution. However, if more information is fused, this condition may not hold.
@@
The SLIM backend remains effective for visual SLAM, as its problem formulation is inherently aligned with the principles of visual bundle adjustment. For instance, in the context of a crowdsourced semantic mapping task on roads, the line and plane features within the framework can be effortlessly replaced with semantic entities such as poles, lane markings, and traffic signs. By incorporating updated landmark definitions and residual formulations, the backend solver retains its adaptability to novel problem representations, thereby ensuring robust and precise solutions.
Despite the focus on urban environments, the SLIM system could also be applied to other structured environments, such as forests and indoors. In forests, the ground and trees can be parameterized into planes and lines. Indoor environments are highly structured with elements like walls and columns, which are typical planes and lines. We also want to emphasize that the goal of the SLIM system is not to achieve state-of-the-art performance. This study primarily aims to provide insights into lightweightness and scalability for LiDAR mapping in urban environments, which are the motivations as stated in Section I.
@@ VIII Conclusion and Future Study
In this study, we present SLIM, a lightweight and scalable LiDAR mapping system for urban environments. Our contributions span from front-end feature extraction to back-end map optimization and management. Experiments conducted on various datasets with different LiDAR sensor types demonstrate the effectiveness of the SLIM system. The results show that SLIM can merge, refine, and maintain new maps using only line and plane representations.
Future studies could explore several promising directions. First, SLIM could be deployed in other structured environments similar to urban scenarios. At the front end, learning-based feature extraction could reduce the need for parameter tuning. At the back end, experiments currently rely on manual region divisions, and a more adaptive approach could be a better choice for long-term autonomy.